OpenStudy (anonymous):

List all the possible rational zeros of p(x)=2x^3-3x^2+4x-6.

OpenStudy (anonymous):

p(x)=2x^3-3x^2+4x-6 y=x^2(2x-3)+2(2x-3) y=(2x-3)(x^2+2) now set it to zero and solve for each of the brackets (idk if we can do this, since it already =y but i can't think of anything else to do...) 0=(2x-3)(x^2+2) 0=2x-3 2x=3 x=3/2 and now the 2nd eqt 0=x^2+2 -2=x^2 root of -2=x (unless you're working with i, this is impossible so we eliminate this answer) so we are left with x=3/2

OpenStudy (mathmale):

You are not being asked to find one actual zero / root / solution of this problem, and there's no equation here, only the definition of a function: p(x)=2x^3-3x^2+4x-6. Please refer to this web page: http://www.virtualnerd.com/algebra-2/polynomials/roots-zeros/rational-zero-theorem/rational-zeros-example and make certain you understand what "rational zeros" are, before trying to solve this problem.

OpenStudy (anonymous):

Okay :)

OpenStudy (mathmale):

Possible rational zeros are identified as follows: List all possible factors of the constant term of your polynomial. In this case that constant term is -6. As an example, I'm typing out for you the factors of -6:\[\pm1,\pm2,\pm3,\pm6\]

OpenStudy (mathmale):

Now do the same thing for the coefficient of your highest-order term. In this problem that coefficient is 2. What are all the possible factors of 2?